OpenStudy (anonymous):

Find the radius of an aluminum cylinder that is 2.00 cm long and has a mass of 12.4 g. V = pi* r^2 * l

hartnn (hartnn):

incomplete information radius is not given mass is useless in finding Volume

OpenStudy (zzr0ck3r):

what is the density of aluminum?

OpenStudy (zzr0ck3r):

v = m/d m/d = pi*r^2*l r = sqrt((m/(d*pi*l))

OpenStudy (anonymous):

Thank you very much @zzr0ck3r lol I never thought about looking that info up.

hartnn (hartnn):

but density is also not given...

OpenStudy (zzr0ck3r):

its known

OpenStudy (zzr0ck3r):

this is prob a physics class not math

OpenStudy (zzr0ck3r):

they say aluminum for a reason:)